### v2

@@ Abstract
The use of ad-hoc engineered viruses in the fight against tumours is one of the greatest ideas in cancer therapeutics within the last three decades. Although some remarkable successes have been obtained, it is still not entirely clear how to achieve reliable protocols that can be routinely employed with confidence on a significant range of tumours. In this work, we thus concentrate on the study of different mathematical descriptions of virotherapy with the aim of better understanding the role of viral infectivity and viral dynamics in positive therapeutic outcomes.
@@ 1 Introduction
Oncolytic virotherapy is an anti-cancer technique that uses genetically modified viruses to directly infiltrate and destroy tumour cells, as well as to elicit an anti-cancer immune response [blanchette23, fountzilas17, Kelly2007651review, lawler17, russell18]. Although it has been known for more than a century that external pathogens can provoke and sustain an immune reaction beneficial to cancer patients, systematic investigations and clinical trials on virotherapy as a viable cure against some tumours have only started a couple of decades ago. Notwithstanding some successes [shalhout23], there are still many unknown and complex factors that have not been completely understood and are preventing a full adoption of virotherapy in the clinics. Furthermore and as a result of the shortcomings for some stand-alone virotherapy approaches, combinations with a number of other therapeutics is also currently being attempted, including, for example, immunotherapies [engeland22], immune checkpoint inhibitors [Sharma2023Checkpoint], chimeric receptor T-cells [Sterner2021CarT] and others  [Mantovani2022Macro, Liu2024Other].
@@
There are many approaches in the mathematical literature that aim at describing the interaction between cancers, viruses and the immune systems. For example, ordinary differential equations (ODEs) have extensively been used [almuallem21, eftimie18, storey20, vithanage23, wodarz01], as well as partial differential equations (PDEs) [friedman18, kim18, lee20, wu04], stochastic agent-based models (ABMs) [storey21, surendran23] and hybrid discrete-continuous multi-scale models [jenner22, paiva09]. The present work is using some of the insights from the work cited above but concentrates on the role of an explicit formulation of the dynamics of viral infection. This aspect represents one of the sensitive points of the approach and its analysis can help improve the general understanding of virotherapy and its efficacy.
@@
In our previous works [morselli23, morselli24], we have modelled infection of tumour cells due to oncolytic viruses without an explicit term for viral particles, i.e. the infection have been implicit and rate-based, since there have been no diffusive and/or motility terms for the viral vectors. We have also proposed two alternative sets of rules for cancer cell movement, based on unhindered or constrained cellular motility. In this work, we instead investigate the effect of an explicit viral term and its impact on outcomes for different movement rules for tumours, with the aim of understanding if and how eradication is governed by viral dynamics. In line with our previous approaches, we employ both an individual-based probabilistic model for single cells’ and viral vectors’ rules and a deterministic continuum model for volume fractions and concentrations. Standard methods are used to derive the continuum macroscopic model from the underlying discrete stochastic model (see, for example, Refs. [chaplain20, champagnat07, johnston15, lorenzimurray20, macfarlane22, penington11]). By comparing the different schemes, stochasticity, oscillations and tumour escape are investigated in the light of the mechanisms for viral infection; the specific focus on viral dynamics is the main novelty of this work. Furthermore, this comparison also allows us to refine the current knowledge on how waves of infections propagate, for small and large cell population numbers, and on how accurately the wave speed can be predicted using an analytic approach.
@@
To highlight why this perspective is important, let us remark that infections may take place via a number of mechanisms, including direct cell-to-cell transmission and cell-free transmission mediated by diffusing virions; the actual combination of processes is hard to establish in full detail (see Ref. [graw16, Cohen2016virusInv] and the references therein). The two mechanisms appear relatively interchangeable in non-spatial models; on the other hand, spatial models may highlight some selected features of the two mechanisms [williams25]. In the specific case of oncolytic virotherapy, there are also several obstacles to viral diffusion in the tumour microenvironment: indeed, factors such as the extracellular matrix composition, immune cell infiltration, tumour stroma and hypoxic regions can impede viral penetration, replication and spread within the tumour [wojton10]. These considerations suggest that viral spread far from infected cells could play a minor role in virotherapy with respect to the infections caused by cell-to-cell contact, membrane fusion, endocytosis and close-range free virions. As a consequence, a common choice has been to neglect the complexities of viral dynamics and only focus on uninfected and infected cells: this approach has been used for non-spatial models of oncolytic viruses [komarova10, novozhilov06] and a few spatial models [morselli23, wodarz12]. On the other hand, the uncertainties highlighted above motivate our interest in exploring how the addition of explicit viral dynamics affects the outcome, at least at a theoretical level. So, we then model the decreased viral displacement by changing the diffusion coefficient, although direct modifications of the equation that directly incorporate obstacles in the diffusion are also possible [pooladvand22]. Our results show that, in most cases, this is comparable to the two-equations model obtained by assuming quasi-stationary viral dynamics. On the other hand, a good viral infiltration of the tumour is associated with a treatment outcome that is at least partially favourable; the mathematical model then provides suggestions on how to improve the outcome.
@@
The paper is organised as follows. In the following Section we derive and discuss the models we use, highlighting in detail the hypotheses and the conditions that are typical of the ABMs, PDEs and ODEs approaches for virotherapy. The subsequent Section contains original results on various typical scenarios, analysing the effects of stochasticity and viral infectivity on eradication, oscillations and tumour escape. Some new results on wave propagation dynamics, analytical evaluation of typical wave speeds and different tumour diffusivity are also of note. Finally, a Conclusion section critically discusses the findings and the limits of the different systems, suggests some insights with possible therapeutic benefits and provides considerations on possible extensions and new avenues for research.
@@ 2.1 Description of the agent-based models
We build upon the modelling framework presented in Ref. [morselli23] and include viral particles, which were not modelled in our previous work: their concentration is described by a discrete non-negative function; as a consequence, the modelling framework is hybrid discrete-continuous. For the sake of simplicity, we restrict our attention to one spatial dimension in the formulation of the models and refer to Remarks 2.1 for the adaptation to two spatial dimensions.
@@
We denote by $\rho_{j}^{n}$ the local pressure at time $t_{n}$ and position $x_{j}$; we assume it to be equal to the local cell density, so that $\rho_{j}^{n}\coloneqq u_{j}^{n}+i_{j}^{n}$. In principle, we could also consider more general expressions for the pressure, such as the functional form proposed in Ref. [perthame14]; while there might be differences in the way the system reaches the carrying capacity and how the model appears in the continuum limit, it has been shown that the overall behaviour of the tumour is not profoundly affected by different functional forms [macfarlane22].
@@
The rules governing the dynamics of the agents correspond precisely to the ones presented in Ref. [morselli23], with the only exception of the infection, which here involves the viral concentration. We consider both movement mechanisms described previously, i.e. undirected and pressure-driven, giving rise to different models. We then consider an additional equation for the evolution of the viral density.
@@
Proliferation, undirected movement and infection of uninfected cancer cells follow the same rules as in Ref. [morselli23], which we recall here briefly for the sake of completeness.
@@
The infection is here caused by contact between uninfected cells and viral particles. This means that an uninfected cell that occupies position $x_{j}$ at time $t_{n}$ becomes infected with probability $\tau{\beta}v_{j}^{n}/K$, where $K$ is the carrying capacity and ${\beta}>0$ is a constant infection rate. Note the different approach with respect to our previous works, in which we assume that an uninfected cell becomes infected upon contact with another infected cell that occupies the same position [morselli23, morselli24]. The difference is the key ingredient to compare cell-to-cell infection with the infection mediated by free virions.
@@
We assume that infected cells release viral particles when they die due to lysis. Since cell death is stochastic, the balance equation for viral density needs to take it into account. The virus also decays at rate $q_{v}>0$ and diffuses. Here, we do consider the uptake of free virions by cancer cells: a few models include this process [bajzer08, pooladvand21], but it is commonly assumed to be negligible. The resulting balance equation is finally:
@@
where $D_{v}>0$ is the diffusion coefficient and $L_{j}^{n}$ denotes the number of infected cells at position $x_{j}$ that die at time $t_{n}$; this last term is multiplied by $\alpha$, which is the number of viral particles released, and divided by $\delta$ to obtain the density from the number of particles. This equation closely resembles the ones used in Refs. [almeida22, bubba20, morselli24] to model the evolution of a chemoattractant concentration, with the main difference given by the presence of stochasticity.
@@ 2.2 Corresponding continuum models
Letting $\tau,\delta\to 0$ in such a way that $\theta\frac{\delta^{2}}{2\tau}\to D$ we arrive at the continuum counterparts of the previously described individual models. In the case of cancer cells, the derivation is almost identical to the calculations performed in Ref. [morselli23], with the only difference that now the term $\tau{\beta}v_{j}^{n}/K$ replaces $\tau{\tilde{\beta}}i_{j}^{n}/K$; for the sake of brevity, we do not report the details. On the other hand, the continuum counterpart of Eq. (2.1) follows the same approach adopted for the chemoattractant in Refs. [almeida22, bubba20, morselli24], which we now sketch: we first take the expected values of Eq. (2.1) and rearrange the terms to obtain
@@
where $\mathbb{E}[\cdot]$ is the expected value. We then assume that there is a function $v\in C^{2}([0,+\infty)\times\mathbb{R})$ such that $v_{j}^{n}=v(t_{n},x_{j})=v$ and let $\tau,\delta\to 0$ to get
@@
$\begin{cases}\partial_{t}u=D\partial_{xx}^{2}u+pu\Bigl(1-\dfrac{u+i}{K}\Bigr)-\dfrac{{\beta}}{K}uv\\[8.0pt] \partial_{t}i=D\partial_{xx}^{2}i+\dfrac{{\beta}}{K}uv-qi\\[8.0pt] \partial_{t}v=D_{v}\partial_{xx}^{2}v+\alpha qi-q_{v}v\end{cases}$ (U3)
Note that this model has already been introduced in Ref. [baabdulla23] (although the derivation from the discrete counterpart, to our knowledge, has not been previously performed) and it is a slightly different version of the approach in Ref. [pooladvand21].
@@
$\begin{cases}\partial_{t}u=\dfrac{D}{K}\partial_{x}[u\partial_{x}(u+i)]+pu\Bigl(1-\dfrac{u+i}{K}\Bigr)-\dfrac{{\beta}}{K}uv\\[8.0pt] \partial_{t}i=\dfrac{D}{K}\partial_{x}[u\partial_{x}(u+i)]+\dfrac{{\beta}}{K}uv-qi\\[8.0pt] \partial_{t}v=D_{v}\partial_{xx}^{2}v+\alpha qi-q_{v}v\end{cases}$ (P3)
This system shares some similarities to the ones presented in Refs. [byrne09, gwiazda19, perthame14], although we are not aware of other works that combine cross-diffusion and standard diffusion in the context of infections. We will see shortly that viral spatial diffusion dampens several of the characteristic features described in the case of pressure-driven movement described previously [morselli23]. It is also interesting to compare Eq. (P3) with Eq. (P2) that will be introduced shortly to highlight the important differences in viral spreading and therapeutic efficiency related to different mechanisms of tumour cell’s movement, especially when viral diffusion is partially constraint.
@@
with $k=u,i$ and $\boldsymbol{e}\in\{(\pm 1,0),(0,\pm 1)\}$ in the case of pressure-driven movement. We then need to scale $\tau$ and $\delta$ in such a way that $\frac{\delta^{2}}{4\tau}\to D$.
@@ 2.3 Quasi-steady viral dynamics in non-spatial models
This choice allows to condense into a single parameter $\tilde{\beta}$ the uncertainties of several biological processes related to the virus, which are not entirely known, and simplify the mathematical analysis (see, for example, Refs. [komarova10, novozhilov06]). Nevertheless, it is important to remark that this system does not exhibit the same oscillatory behaviours of Eq. (2.2) and this may lead to overlook some relevant situations, as we discuss in Subsection 3.4.
@@
$(u^{*},i^{*})\coloneqq\Bigl(\frac{qK}{\tilde{\beta}},\frac{pK(\tilde{\beta}-q)}{\tilde{\beta}(\tilde{\beta}+p)}\Bigr)$ (2.6)
The first equilibrium point is always unstable. The second one is stable when $\tilde{\beta}<q$ (i.e., $i^{*}<0$) and unstable when $\tilde{\beta}>q$ (i.e., $i^{*}>0$); in this last case, the third equilibrium is stable (see also Ref. [morselli23]).
@@
$\Bigl(u^{*},i^{*},\frac{\alpha q}{q_{v}}\,i^{*}\Bigr)=\Bigl(\frac{qK}{\tilde{\beta}},\frac{pK(\tilde{\beta}-q)}{\tilde{\beta}(\tilde{\beta}+p)},\frac{\alpha q}{q_{v}}\frac{pK(\tilde{\beta}-q)}{\tilde{\beta}(\tilde{\beta}+p)}\Bigr)$ (2.7)
and it may become unstable for $\tilde{\beta}>q$ when a stable limit cycle arises (as described in Ref. [baabdulla23]). Let us remark that the spatially homogeneous steady-state of the tumour depends on $\alpha$ and $q$ only through $\tilde{\beta}$, hence we should expect that different parameter combinations that maintain the ratio $\alpha/q_{v}$ unchanged result in similar dynamics. This situation is analysed in Subsection 3.2.

### v3

@@ Abstract
The use of ad-hoc engineered viruses in the fight against tumours is one of the greatest ideas in cancer therapeutics within the last three decades. Although some remarkable successes have been obtained, it is still not entirely clear how to achieve reliable protocols that can be routinely employed with confidence on a significant range of tumours. In this work, we concentrate on the study of different mathematical descriptions of virotherapy with the aim of better understanding the role of viral infectivity and viral dynamics in positive therapeutic outcomes.
@@ 1 Introduction
Oncolytic virotherapy is an anti-cancer technique that uses genetically modified viruses to directly infiltrate and destroy tumour cells, as well as to elicit an anti-cancer immune response [7, 25, 35, 40, 65]. Although it has been known for more than a century that external pathogens can provoke and sustain an immune reaction beneficial to cancer patients, systematic investigations and clinical trials on virotherapy as a viable cure against some tumours have only started a couple of decades ago. Notwithstanding some successes [66], there are still many unknown and complex factors that have not been completely understood and are preventing a full adoption of virotherapy in the clinics. Furthermore and as a result of the shortcomings for some stand-alone virotherapy approaches [50], combinations with a number of other therapeutics is also currently being attempted, including, for example, immunotherapies [23], immune checkpoint inhibitors [67], chimeric receptor T-cells [68] and others  [49, 44].
@@
There are many approaches in the mathematical literature that aim at describing the interaction between cancers, viruses and the immune systems. For example, ordinary differential equations (ODEs) have extensively been used [2, 22, 70, 72, 76], as well as partial differential equations (PDEs) [26, 37, 41, 80], stochastic agent-based models (ABMs) [69, 71] and hybrid discrete-continuous multi-scale models [32, 57]. The present work uses some of the insights from previous works, focusing on the role of an explicit formulation of the dynamics of viral infection. This aspect represents one of the sensitive points of the approach and its analysis can help improve the general understanding of virotherapy.
@@
In our previous works [53, 54], we have modelled infection of tumour cells due to oncolytic viruses without an explicit term for viral particles, i.e. the infection have been implicit and rate-based, since there have been no diffusive and/or motility terms for the viral vectors. We have also proposed two alternative sets of rules for cancer cell movement, based on unhindered or constrained cellular motility. In this work, we instead investigate the effect of an explicit viral term and its impact on outcomes for different movement rules for tumours, with the aim of understanding if and how eradication is governed by viral dynamics. In line with our previous approaches, we employ both an individual-based probabilistic model for single cells’ and viral vectors’ rules and a deterministic continuum model for volume fractions and concentrations. Standard methods are used to derive the continuum macroscopic model from the underlying discrete stochastic model (see, for example, Refs. [14, 13, 33, 47, 48, 58]). By comparing the different schemes, stochasticity, oscillations and tumour escape are investigated in the light of the mechanisms for viral infection; the specific focus on viral dynamics is the main novelty of this work. Furthermore, this comparison also allows us to refine the current knowledge on how waves of infections propagate, for small and large cell population numbers, and on how accurately the wave speed can be predicted using an analytic approach.
@@
To highlight why this perspective is important, let us remark that infections may take place via a number of mechanisms, including direct cell-to-cell transmission and cell-free transmission mediated by diffusing virions; the actual combination of processes is hard to establish in full detail (see Refs. [30, 17] and the references therein). The two mechanisms appear relatively interchangeable in non-spatial models; on the other hand, spatial models may highlight some selected features of the two mechanisms [75]. In the specific case of oncolytic virotherapy, there are also several obstacles to viral diffusion in the tumour microenvironment: indeed, factors such as the extracellular matrix composition, immune cell infiltration, tumour stroma and hypoxic regions can impede viral penetration, replication and spread within the tumour [78]. These considerations suggest that viral spread far from infected cells could play a minor role in virotherapy with respect to the infections caused by cell-to-cell contact, membrane fusion, endocytosis and close-range free virions. As a consequence, a common choice has been to neglect the complexities of viral dynamics and only focus on uninfected and infected cells: this approach has been used for non-spatial models of oncolytic viruses [39, 56] and a few spatial models [53, 77]. On the other hand, the uncertainties highlighted above motivate our interest in exploring how the addition of explicit viral dynamics affects the outcome, at least at a theoretical level. We then model the decreased viral displacement by changing the diffusion coefficient, although direct modifications of the equation that directly incorporate obstacles in the diffusion are also possible [62]. Our results show that, in most cases, this is comparable to the two-equations model obtained by assuming quasi-stationary viral dynamics. On the other hand, a good viral infiltration of the tumour is associated with a treatment outcome that is at least partially favourable; the mathematical model then provides suggestions on how to improve the outcome.
@@
The paper is organised as follows. In Section 2, we derive and discuss the models we use, highlighting in detail the hypotheses and the conditions that are typical of the ABMs, PDEs and ODEs approaches for virotherapy. Section 3 contains original results on various typical scenarios, analysing the effects of stochasticity and viral infectivity on eradication, oscillations and tumour escape. Some new results on wave propagation dynamics, analytical evaluation of typical wave speeds and different tumour diffusivity are also of note. Finally, Section 4 critically discusses the findings and the limits of the different systems, suggests some insights with possible therapeutic benefits and provides considerations on possible extensions and new avenues for research.
@@ 2.1 Description of the agent-based models
We build upon the modelling framework presented in Ref. [53] and include viral particles, which were not modelled in our previous work: their concentration is described by a discrete non-negative function; as a consequence, the modelling framework is hybrid discrete-continuous. For the sake of simplicity, we restrict our attention to one spatial dimension in the formulation of the models and refer to Remark 2.1 for the adaptation to two spatial dimensions.
@@
We denote by $\rho_{j}^{n}$ the local pressure at time $t_{n}$ and position $x_{j}$; we assume it to be equal to the local cell density, so that $\rho_{j}^{n}\coloneqq u_{j}^{n}+i_{j}^{n}$. In principle, we could also consider more general expressions for the pressure, such as the functional form proposed in Ref. [59]; while there might be differences in the way the system reaches the carrying capacity and how the model appears in the continuum limit, it has been shown that the overall behaviour of the tumour is not profoundly affected by different functional forms [48].
@@
The rules governing the dynamics of the agents correspond precisely to the ones presented in Ref. [53], with the only exception of the infection, which here involves the viral concentration. We consider both movement mechanisms described previously, i.e. undirected and pressure-driven, giving rise to different models. We then consider an additional equation for the evolution of the viral density.
@@
Proliferation, undirected movement and infection of uninfected cancer cells follow the same rules as in Ref. [53], which we recall here briefly for the sake of completeness.
@@
We also assume that at every time step an infected cell may die because of lysis with probability $\tau q$, where $q>0$ is a constant death rate. The infection is here caused by contact between uninfected cells and viral particles. This means that an uninfected cell that occupies position $x_{j}$ at time $t_{n}$ becomes infected with probability $\tau{\beta}v_{j}^{n}/K$, where $K$ is the carrying capacity and ${\beta}>0$ is a constant infection rate. Note the different approach with respect to our previous works, in which we assume that an uninfected cell becomes infected upon contact with another infected cell that occupies the same position [53, 54]. This difference is the key ingredient to compare cell-to-cell infection with the infection mediated by free virions.
@@
We assume that infected cells release viral particles when they die due to lysis. Since cell death is stochastic, the balance equation for viral density needs to take it into account. The virus also decays at rate $q_{v}>0$ and diffuses. We use the notation $q_{v}$ so that parameters related to decays are labelled in a consistent way. Here, we do not consider the uptake of free virions by cancer cells: a few models include this process [5, 63], but it is commonly assumed to be negligible. The resulting balance equation is finally:
@@
where $D_{v}>0$ is the diffusion coefficient and $L_{j}^{n}$ denotes the number of infected cells at position $x_{j}$ that die at time $t_{n}$; this last term is multiplied by $\alpha$, which is the number of viral particles released, and divided by $\delta$ to obtain the density from the number of particles. This equation closely resembles the ones used in Refs. [1, 9, 54] to model the evolution of a chemoattractant concentration, with the main difference given by the presence of stochasticity.
@@ 2.2 Corresponding continuum models
Letting $\tau,\delta\to 0$ in such a way that $\theta\frac{\delta^{2}}{2\tau}\to D_{k}$ (with $k=U,P$ in the two different models), we arrive at the continuum counterparts of the individual models described above. The values of the diffusion coefficients are chosen to ensure identical propagation speeds in both models (see Table 1 and Appendix A.1 for details).
In the case of cancer cells, the derivation is almost identical to the calculations performed in Ref. [53], with the only difference that now the term $\tau{\beta}v_{j}^{n}/K$ replaces $\tau{\tilde{\beta}}i_{j}^{n}/K$; for the sake of brevity, we do not report the details. On the other hand, the continuum counterpart of Eq. (2.1) follows the same approach adopted for the chemoattractant in Refs. [1, 9, 54], which we now sketch: we first take the expected values of Eq. (2.1) and rearrange the terms to obtain
@@
where $\mathbb{E}[\cdot]$ is the expected value. We then assume that there is a function $v\in C^{2}([0,+\infty)\times\mathbb{R})$ such that $v_{j}^{n}=v(t_{n},x_{j})=v$ and let $\tau,\delta\to 0$ to obtain
@@
$\begin{cases}\partial_{t}u=D_{U}\partial_{xx}^{2}u+pu\Bigl(1-\dfrac{u+i}{K}\Bigr)-\dfrac{{\beta}}{K}uv\\[8.0pt] \partial_{t}i=D_{U}\partial_{xx}^{2}i+\dfrac{{\beta}}{K}uv-qi\\[8.0pt] \partial_{t}v=D_{v}\partial_{xx}^{2}v+\alpha qi-q_{v}v\end{cases}$ (U3)
Note that this model has already been introduced in Ref. [4] (although the derivation from the discrete counterpart, to our knowledge, has not been previously performed) and it is a slightly different version of the approach in Ref. [63].
@@
$\begin{cases}\partial_{t}u=\dfrac{D_{P}}{K}\partial_{x}[u\partial_{x}(u+i)]+pu\Bigl(1-\dfrac{u+i}{K}\Bigr)-\dfrac{{\beta}}{K}uv\\[8.0pt] \partial_{t}i=\dfrac{D_{P}}{K}\partial_{x}[i\partial_{x}(u+i)]+\dfrac{{\beta}}{K}uv-qi\\[8.0pt] \partial_{t}v=D_{v}\partial_{xx}^{2}v+\alpha qi-q_{v}v\end{cases}$ (P3)
This system shares some similarities with the ones presented in Refs. [11, 31, 59], although we are not aware of other works that combine cross-diffusion and standard diffusion in the context of infections. The well-posedness of similar kinds of cross-diffusion “triangular” systems (i.e., with some equations only containing standard diffusion) have been previously studied [8, 18, 19, 20, 81], but to our knowledge the existing results cannot be directly applied to Eq. (P3). We will see shortly that viral spatial diffusion dampens several of the characteristic features described in the case of pressure-driven movement studied previously [53]. It is also interesting to compare Eq. (P3) with Eq. (P2) that will be introduced shortly to highlight the important differences in viral spreading and therapeutic efficiency related to different mechanisms of tumour cell’s movement, especially when viral diffusion is partially constrained.
@@
with $k=u,i$ and $\boldsymbol{e}\in\{(\pm 1,0),(0,\pm 1)\}$ in the case of pressure-driven movement. We then need to scale $\tau$ and $\delta$ in such a way that $\frac{\delta^{2}}{4\tau}\to D_{k}$, with $k=U,P$ in the two different models.
@@ 2.3 Quasi-steady viral dynamics in non-spatial models
This choice allows to condense into a single parameter $\tilde{\beta}$ the uncertainties of several biological processes related to the virus, which are not entirely known, and simplify the mathematical analysis (see, for example, Refs. [39, 56]). Nevertheless, it is important to remark that this system does not exhibit the same oscillatory behaviours of Eq. (2.2) and this may lead to overlook some relevant situations, as we discuss in Subsection 3.4.
@@
$(u^{*},i^{*})\coloneqq\left(\frac{qK}{\tilde{\beta}},\frac{pK(\tilde{\beta}-q)}{\tilde{\beta}(\tilde{\beta}+p)}\right)$ (2.6)
The first equilibrium point is always unstable. The second one is stable when $\tilde{\beta}<q$ (i.e., $i^{*}<0$) and unstable when $\tilde{\beta}>q$ (i.e., $i^{*}>0$); in this last case, the third equilibrium is stable (see also Ref. [53]).
@@
$\Bigl(u^{*},i^{*},\frac{\alpha q}{q_{v}}\,i^{*}\Bigr)=\left(\frac{qK}{\tilde{\beta}},\frac{pK(\tilde{\beta}-q)}{\tilde{\beta}(\tilde{\beta}+p)},\frac{\alpha q}{q_{v}}\frac{pK(\tilde{\beta}-q)}{\tilde{\beta}(\tilde{\beta}+p)}\right)$ (2.7)
and it may become unstable for $\tilde{\beta}>q$ when a stable limit cycle arises (as described in Refs. [4, 6]). Let us remark that the spatially homogeneous steady-state of the tumour depends on $\alpha$ and $q_{v}$ only through $\tilde{\beta}$, hence we should expect similar dynamics for parameter combinations that maintain the ratio $\alpha/q_{v}$ unchanged. This situation is analysed in Subsection 3.2.